OpenStudy (thickgirl_angel):

Find the 4th term if the sequend in which a1 = 2 and an+1 = -4an + 2.

OpenStudy (sshayer):

put n=1,find a2 put n=2 ,find a3 put n=3,find a4

OpenStudy (sshayer):

correction a2=-4a1+2=-4*2+2=-8+2=-6